# Constrained Discrete Diffusion

### v2

@@ Abstract
Discrete diffusion models are a class of generative models that construct sequences by progressively denoising samples from a categorical noise distribution. Beyond their rapidly growing ability to generate coherent natural language, these models present a new and important opportunity to enforce sequence-level constraints, a capability that current autoregressive models cannot natively provide. This paper capitalizes on this opportunity by introducing Constrained Discrete Diffusion (CDD), a novel integration of differentiable constraint optimization within the diffusion process to ensure adherence to constraints, logic rules, or safety requirements for generated sequences. Unlike conventional text generators that often rely on post-hoc filtering or model retraining for controllable generation, CDD directly imposes constraints into the discrete diffusion sampling process, resulting in a training-free and effective approach. Experiments in toxicity-controlled text generation, property-constrained molecule design, and instruction-constrained text completion demonstrate that CDD achieves zero constraint violations in a diverse array of tasks while preserving fluency, novelty, and coherence while outperforming autoregressive and existing discrete diffusion approaches.
@@ 1 Introduction
Language generation, alongside many scientific problems, admits a natural representation as the generation of a discrete sequence from a finite alphabet. Examples range from natural language to molecular SMILES strings and linearized chemical procedures [1, 2, 3, 4]. While large language models (LLMs) and related sequence transformers have recently accelerated discovery by proposing candidate sequences with desirable attributes, their autoregressive sampling mechanism produces tokens sequentially, hindering the ability to provide a native mechanism to ensure constrained feasibility. When these constraints are not satisfied, the generated outputs may be unreliable and ineffective in real-world applications. For example, a single out-of-vocabulary atom in a SMILES string can render a synthesized compound meaningless or even dangerous; similarly an overlooked volume unit in an autonomous laboratory protocol can trigger unsafe reactions.
To limit such risks, LLMs are often deployed with a variety of guardrails. These include soft alignment through reinforcement learning from human feedback (RLHF), rejection sampling, or heuristic post-processing [5, 3]. However, these methods do not offer provable compliance with logical formulas, stoichiometric rules, or regulatory thresholds. This is exacerbated by the autoregressive nature of common language models, which generate sequences one token at a time, making it difficult to enforce constraints at the sequence level.
Figure 1: Comparison of Constrained Discrete Diffusion (CDD) and baseline models. CDD imposes constraints without sacrificing fluency or expressiveness.
In contrast, discrete diffusion models offer a compelling alternative generative mechanism [6, 7, 8, 9]. They refine a fully corrupted sequence by iteratively denoising the entire sample, and have shown strong performance across a variety of tasks, including text generation, molecular design, and program synthesis [4, 10, 7]. Since each step exposes a global view of the partial sequence, it creates a natural opportunity to impose sequence-level structure. This work introduces Constrained Discrete Diffusion (CDD), a framework that capitalizes on this property by coupling discrete diffusion with a differentiable projection operator. Given a corrupted sequence at a particular diffusion step, the projection searches for a new candidate that stays close to the model’s current score distribution, preserves entropy-based diversity, and simultaneously satisfies all user-defined constraints before the next denoising update. Because the introduced projection acts only at sampling time, the method is training-free and needs neither model retraining nor post hoc filtering. Its stochastic formulation also preserves the generative diversity, which is crucial for exploratory scientific design, as will be shown in the empirical analysis.
@@ 2 Related Work
Recent efforts in controllable text generation have largely focused on constrained decoding, where the output is dynamically restricted to adhere to predetermined syntactic, semantic, or lexical rules. Approaches in this area modify the decoding process to prune the vocabulary, ensuring that only tokens compliant with a formal grammar are considered [11], or employ external modules to filter outputs when direct access to the model’s logits is limited [12]. Other methods have also adapted beam search to encourage the presence or absence of specific words [13, 14]. Although these techniques effectively guide the generation process, they depend on augmented sampling heuristics which encourage, but frequently fail to provide, satisfaction of even simple constraint sets [15]. To address this, other research has implemented mappings from natural language to formal logic allowing verification of factual and logical consistency [16, 17, 18] and enforcement of feasibility constraints with external modules [19, 20]. However, these approaches, while effective in their respective domains, often require additional, domain-specific components to ensure that the generated outputs are not only syntactically valid but also adhere to practical or logical requirements.
@@
Other controllable generation approaches rely on energy-based sampling procedures to impose constraints on the outputs. These methods typically define an energy function combining language model likelihood with constraint penalties and leverage algorithms such as Metropolis-Hastings to sample low-energy (i.e., high probability) sequences [21, 22, 23]. While such methods are highly effective for encouraging simple properties (e.g., style transfer), they inherit many of the theoretical and practical shortcomings of rejection sampling. In particular, hard constraints are typically enforced only probabilistically, making energy-based approaches less suitable for settings requiring strict constraint satisfaction, as is desirable for the tasks considered in this work.
@@
Finally, several gradient-based sampling frameworks have been proposed to impose constraints on token sequences [24]. Building on this approach, Amini et al. [25] improve generation quality via structured Hamiltonian Monte Carlo. However, these methods lack mechanisms for enforcing hard constraints, and are limited to soft attribute control. Other gradient-based control methods for autoregressive generation, such as PPLM [26], apply conditional steering during decoding to bias generations toward desired attributes; we compare against this methods directly, and show that while it can shift output distributions, it fails to reliably enforce structural or semantic constraints.
3 Preliminaries
Generative Diffusion Models.
Generative diffusion models [27, 28] have revolutionized high-fidelity creation of complex data and serve as key components in state-of-the-art image and video synthesis pipelines [29, 30, 31]. Formally, they operate in two phases: a forward phase, which gradually introduces noise to clean data samples, followed by a reverse denoising phase, in which a deep neural network is trained to iteratively remove noise. The forward process defines a Markov chain $\{\bm{x}_{t}\}_{t=0}^{T}$, with initial sample $\bm{x}_{0}\sim p(\bm{x}_{0})$ and each transition $q(\bm{x}_{t}|\bm{x}_{s})$ obtained by adding Gaussian noise that increases as $t\to T$. Here, $s$ is the subsequent timestep in the continuous time diffusion process and $0\leq s<t\leq T$. The reverse process relies on a learned network that directly predicts the noise added at each step, enabling the gradual recovery of the original data distribution. Once trained, this network can then be used to iteratively “denoise” random noise samples back into data-like samples. This step is also known as sampling.
Discrete Diffusion Language Models.
@@ 3 Preliminaries
While diffusion models were originally developed for continuous data, they have recently been extended to discrete domains such as text, enabling non-autoregressive generation [32, 9, 7, 33]. In contrast to autoregressive models which predict tokens one by one, discrete diffusion methods generate entire sequences in parallel by first corrupting sequences through a forward noising process and then iteratively reconstructing them with a learned reverse process. This is a key enabler recognized by this work, which exploits this modus operandi to impose global constraints while simultaneously maintaining high fidelity.
@@
where $\alpha_{t}$ is a schedule that decreases over time, $\mathrm{Cat}(\cdot;p)$ is the categorical distribution over probability vector $p\in\Delta^{N}$, and $\nu$ is a fixed reference distribution that specifies the type of corruption applied to each token. For example, the Masked Diffusion Language Model (MDLM) [7] uses a one-hot vector corresponding to the [MASK] token for $\nu$. In contrast, Uniform Diffusion Language Models (UDLM) [8], uses $\nu$ as the uniform distribution over the vocabulary. Conversely, this instantiation allows tokens to be re-perturbed in later time steps.
@@
$q(\bm{x}_{s}\mid\bm{x}_{t},\bm{x}_{0})=\begin{cases}\mathrm{Cat}\bigl{(}\bm{x}_{s};\bm{x}_{t}\bigr{)},&\text{if }\bm{x}_{t}\neq\bm{\nu},\\[4.0pt] \mathrm{Cat}\!\Bigl{(}\bm{x}_{s};\frac{(1-\alpha_{s})\,\bm{\nu}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{\theta}(\bm{x}_{t},t)}{1-\alpha_{t}}\,\Bigr{)},&\text{if }\bm{x}_{t}=\bm{\nu}.\end{cases}$ (2)
Equation (2) formalizes this reverse process parameterization, where $\mathbf{x}_{\theta}(\bm{x}_{t},t)$ denotes the trained denoiser.
4 Constraint Discrete Diffusion (CDD)
Projected diffusion sampling.
We begin by introducing a perspective of the discrete diffusion reverse process that motivates our approach. Prior work has shown that continuous score-based diffusion sampling processes can be framed as an sequential optimization procedure [34]. Score-based parameterizations enable this framing as the model learns to predict the gradients of the data density function, $\mathbf{x_{\theta}}(\bm{x}_{t},t)\approx\nabla\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})$ through Score Matching, and these gradients can then be applied to optimize $\bm{x}_{t}$ to the data distribution $q_{t}(\bm{x}_{t})$. This is typically conducted through a Langevin dynamics algorithms [35], which is used either directly [36] or through predictor-corrector Euler-discretizations [37]. In both cases, a series of $M$ steps of Langevin dynamics are taken for a fixed distribution $q_{t}(\bm{x}_{t})$:
@@ 4 Constraint Discrete Diffusion (CDD)
Discrete diffusion models can leverage a discrete generalization of the score function, referred to as the Concrete score, to approximate the gradient of the probability density function [7, 9, 38]. Concrete Score Matching provides an approach mirroring continuous Score Matching in that the estimated gradients of the probability density function are used to guide the sample to high density regions of the target distribution. While not always explicitly framed as Concrete Score Matching [9, 38], the denoiser often implicitly models the score function, as supported by theoretical results in [7] demonstrating its equivalence in simplified formulations ($\nabla\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})\approx\langle\mathbf{x}_{\theta}(\bm{x}_{t},t),\mathbf{y}\rangle\;\;\forall\mathbf{y}=1,\ldots,N$, in this case). This enables the use of Langevin-based samplers for discrete diffusion, as commonly employed, e.g., in [38].
Effectively, under some regularity conditions, Langevin dynamics will converge towards a stationary point. As shown by Xu et al., Langevin dynamics acts as an “almost-minimizer” on the optimized function, which in this case will be the negative probability density function, converging within a fixed bound of the optimum. Hence, for each series of steps with a stationary density function, within this fixed bound, the sampling procedure can be framed as the optimization problem,
@@
Here, Equation (4) extends the representation from an unconstrained optimization problem to a constrained version by introducing a constraint set $\mathbf{C}$. However, while iteratively applying the denoiser enables sampling from the posterior distribution, it does not natively incorporate externally defined constraints, or even implicit ones, given the stochastic nature of the process. Previous work for continuous modalities has proposed enforcing $\bm{x}_{t}\in\mathbf{C}$ by applying a Euclidean projection after each denoising step [34], which is natural for continuous modalities which fall in a real space, but this is misaligned when applied to discrete diffusion which operates over a probability simplex $\bm{x}_{t}\in\Delta^{N}$.
@@
Importantly, while state transitions are imposed on the token probability distributions, constraint satisfaction is evaluated on the decoded sequence. Indeed, the constraints are formulated such that the $\arg\max$ of the projected sequence $\bm{y}$ (referred to as $\bm{y}^{\star}$), must satisfy sentence level criteria $\bm{y}^{\star}\in\mathbf{C}$ [^1] We assume a greedy decoding scheme as is standard to current diffusion-based language models. However, this use of a non-differentiable $\arg\max$ operation also poses challenges to the resolution of the projection, which relies on gradient-based optimization.
Differentiable projection. To address the non-differentiability of the $\arg\max$ function, we apply a Gumbel-Softmax relaxation $\tilde{\phi}$, which approximates discrete token probabilities with continuous values [40] as
@@
where $\bm{x}_{t}^{i}(v)$ is the probability of token $v$ in the vocabulary, $\xi_{v}$ is drawn from the Gumbel$(0,1)^{1}$ distribution for token $v$, and $\text{T}_{\text{sample}}>0$ is the temperature parameter controlling the smoothness of the output. This enables gradient propagation during the projection step, while closely approximating the discrete $\arg\max$ operation.
Augmented Lagrangian projection.
@@
In practice, $\Delta g$ is non-linear, and, thus, to implement Equation (5), we adopt an augmented Lagrangian approach [41]. In augmented Lagrangian methods, the problem constraints are incorporated into the objective of a minimizer via Lagrange multipliers $\lambda$ and a quadratic penalty term $\mu$. Let $\bm{x}_{t}^{\prime}$ be the probability distribution after the denoising step at diffusion time $t$. We introduce a projected distribution $\bm{y}$, which is iteratively updated to reduce the constraint violations (measured by the score $\Delta g$) while remaining close (in KL-divergence) to $\bm{x}_{t}^{\prime}$. Concretely the augmented Lagrangian dual function is defined as:
@@
and is a lower bound of the original projection operator (5) by weak duality [41]. To obtain the strongest Lagrangian relaxation of the projection, the Lagrangian dual can be used to find the best Lagrangian multipliers $\lambda$ and penalty terms $\mu$, i.e.,
@@
Algorithm 1 Augmented Lagrangian Method
1:Init: $\lambda,\mu,\eta,\alpha,\delta$
2:Input: probability distribution $\bm{x}_{t}^{\prime}$
3:$\bm{y}=\bm{x}_{t}^{\prime}$
4:while $\Delta g(\bm{y}^{*})<\delta$ do
5:     for $j\leftarrow 1$ to max_inner_iter do
6:         $\mathcal{L}_{KL}\leftarrow D_{\mathrm{KL}}(\bm{x}_{t}^{\prime}\|\bm{y})$
7:         $\mathcal{L}_{\text{viol}}\leftarrow\sum_{i=1}^{m}\lambda_{i}\,\Delta g_{i}(\tilde{\phi}(\bm{y}))+\tfrac{\mu_{i}}{2}\,\Delta g(\tilde{\phi}(\bm{y}))^{2}$
8:         $\mathcal{L}_{\text{ALM}}\leftarrow\mathcal{L}_{KL}+\mathcal{L}_{\text{viol}}$
9:         ${\bm{y}}\leftarrow{\bm{y}}-\eta\,\nabla_{\bm{y}}\mathcal{L}_{\text{ALM}}$
10:     end for
11:     $\lambda\leftarrow\lambda+\mu\,\Delta g(\bm{y}^{*})$
12:     $\mu\leftarrow\min\bigl{(}\alpha\mu,\,\mu_{\text{max}}\bigr{)}$
13:end while
14:$\bm{x}_{s}=\bm{y}$
15:return $\bm{x}_{s}$
The optimization of (7) proceeds iteratively, following a gradient-based update on $\bm{y}$ while dynamically adjusting the Lagrange multiplier $\lambda$ and penalty coefficient $\mu$. Specifically, we perform the following updates [42]:
@@
To assess the robustness of this projection formulation, we performed an ablation study (Appendix D) examining the sensitivity of the Lagrangian hyper-parameters. Across the full grid of settings, the Lagrangian relaxation converges to feasible solutions in all our test cases.
@@
Let $\bm{C}$ be non-empty and $\beta$-prox-regular in the sense of [43, Def. 13.27], and the score network satisfy $\|\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})\|\leq G$ (a standard consequence of the bounded-data domain after normalization). Then, for positive step sizes $\gamma_{t},\leq\frac{1}{2G^{2}}\beta$, the following inequality holds for the distance to the feasible set $\bm{C}$:
@@
Theorem 4.1 shows that the distance to the feasible set $\bm{C}$ decreases at a rate of $1-\bm{\upalpha}_{t}$ at each step (up to an additive $\bm{\upalpha}_{t}^{2}G^{2}$ noise). This implies $\varepsilon$-feasibility after $\mathcal{O}(\bm{\upalpha}_{\min}^{-1})$ steps, with $\bm{\upalpha}_{\min}\propto\min_{t}\gamma_{t}$, and a cumulative KL drift within $\mathcal{O}(\sum_{t}\bm{\upalpha}_{t})$. Missing proofs are provided in Appendix F.

### v3

@@ Abstract
Discrete diffusion models are a class of generative models that construct sequences by progressively denoising samples from a categorical noise distribution. Beyond their rapidly growing ability to generate coherent natural language, these models present a new and important opportunity to enforce sequence-level constraints, a capability that current autoregressive models cannot natively provide. This paper capitalizes on this opportunity by introducing Constrained Discrete Diffusion (CDD), a novel integration of differentiable constraint optimization within the diffusion process to ensure adherence to constraints, logic rules, or safety requirements for generated sequences. Unlike conventional text generators that often rely on post-hoc filtering or model retraining for controllable generation, CDD directly imposes constraints into the discrete diffusion sampling process, resulting in a training-free and effective approach. Experiments in toxicity-controlled text generation, property-constrained molecule design, and instruction-constrained text completion demonstrate that CDD achieves zero constraint violations in a diverse array of tasks while preserving fluency, novelty, and coherence, while outperforming autoregressive and existing discrete diffusion approaches.
@@ 1 Introduction
Language generation, alongside many scientific problems, admits a natural representation as the generation of a discrete sequence from a finite alphabet. Examples range from natural language to molecular SMILES strings and linearized chemical procedures [li2024embodied, jia2024decision, zhai2024fine, ye2024beyond]. While large language models (LLMs) and related sequence transformers have recently accelerated discovery by proposing candidate sequences with desirable attributes, their autoregressive sampling mechanism produces tokens sequentially, hindering the ability to provide a native mechanism to ensure constrained feasibility. When these constraints are not satisfied, the generated outputs may be unreliable and ineffective in real-world applications. For example, a single out-of-vocabulary atom in a SMILES string can render a synthesized compound meaningless or even dangerous; similarly an overlooked volume unit in an autonomous laboratory protocol can trigger unsafe reactions.
To limit such risks, LLMs are often deployed with a variety of guardrails. These include soft alignment through reinforcement learning from human feedback (RLHF), rejection sampling, or heuristic post-processing [ouyang2022training, zhai2024fine]. However, these methods do not offer provable compliance with logical formulas, stoichiometric rules, or regulatory thresholds. This is exacerbated by the autoregressive nature of
Figure 1: Comparison of Constrained Discrete Diffusion and baseline models. CDD imposes constraints without sacrificing fluency or expressiveness.
common language models, which generate sequences one token at a time, making it difficult to enforce constraints at the sequence level.
In contrast, discrete diffusion models offer a compelling alternative generative mechanism [austin2023structureddenoisingdiffusionmodels, sahoo2024simple, schiff2024simple, loudiscrete]. They refine a fully corrupted sequence by iteratively denoising the entire sample, and have shown strong performance across a variety of tasks, including text generation, molecular design, and program synthesis [ye2024beyond, zheng2024masked, sahoo2024simple]. Since each step exposes a global view of the sequence, it creates a natural opportunity to impose sequence-level structure. This work introduces Constrained Discrete Diffusion (CDD), a framework that capitalizes on this property by coupling discrete diffusion with a differentiable projection operator. Given a corrupted sequence at a particular diffusion step, the projection searches for a new candidate that stays close to the model’s current score distribution, preserves entropy-based diversity, and simultaneously satisfies all user-defined constraints before the next denoising update. Because the introduced projection acts only at sampling time, the method is training-free and needs neither model retraining nor post hoc filtering. Its stochastic formulation also preserves the generative diversity, which is crucial for exploratory scientific design, as will be shown in the empirical analysis.
@@ 2 Related Work
Recent efforts in controllable text generation have largely focused on constrained decoding, where the output is dynamically restricted to adhere to predetermined syntactic, semantic, or lexical rules. Approaches in this area modify the decoding process to prune the vocabulary, ensuring that only tokens compliant with a formal grammar are considered [geng-etal-2023-grammar], or employ external modules to filter outputs when direct access to the model’s logits is limited [geng2024sketch]. Other methods have also adapted beam search to encourage the presence or absence of specific words [lu2020neurologic, lu2021neurologic]. Although these techniques effectively guide the generation process, they depend on augmented sampling heuristics which encourage, but frequently fail to provide, satisfaction of even simple constraint sets [Yang_2021]. To address this, other research has implemented mappings from natural language to formal logic allowing verification of factual and logical consistency [gopalan2018sequence, liu2023grounding, liu2024logic] and enforcement of feasibility constraints with external modules [chen2023autotamp, lin2023text2motion]. However, these approaches, while effective in their respective domains, often require additional, domain-specific components to ensure that the generated outputs are not only syntactically valid but also adhere to practical or logical requirements.
@@
Other controllable generation approaches rely on energy-based sampling procedures to impose constraints on the outputs. These methods typically define an energy function combining language model likelihood with constraint penalties and leverage algorithms such as Metropolis-Hastings to sample low-energy (i.e., high probability) sequences [miao2019cgmh, forristal2023block, gareev2024local]. While such methods are highly effective for encouraging simple properties (e.g., style transfer), they inherit many of the theoretical and practical shortcomings of rejection sampling. In particular, hard constraints are typically enforced only probabilistically, making energy-based approaches less suitable for settings requiring strict constraint satisfaction, as is desirable for the tasks considered in this work.
@@
Finally, several gradient-based sampling frameworks have been proposed to impose constraints on token sequences [kumar2022gradient]. Building on this approach, amini2023structured improve generation quality via structured Hamiltonian Monte Carlo. However, these methods lack mechanisms for enforcing hard constraints, and are limited to soft attribute control. Other gradient-based control methods for autoregressive generation, such as PPLM [dathathri2019plug], apply conditional steering during decoding to bias generations toward desired attributes; we compare against this methods directly, and show that while it can shift output distributions, it fails to reliably enforce structural or semantic constraints. Guided discrete diffusion methods have also been explored for protein design. For example, [gruver2023proteindesignguideddiscrete] applies gradient-based guidance to the denoiser’s hidden states during sampling to steer sequences toward desired property objectives. Compared to these gradient-based guidance methods, our method frames each denoising step as a constrained optimization and enforces feasibility via an augmented-Lagrangian projection, presenting different behavior and formal guarantees.
3 Preliminaries: Discrete Diffusion Models
@@
While diffusion models were originally developed for continuous data, they have recently been extended to discrete domains such as text, enabling non-autoregressive generation [austin2021structured, loudiscrete, sahoo2024simple, shi2024simplified]. In contrast to autoregressive models which predict tokens one by one, discrete diffusion methods generate entire sequences in parallel by first corrupting sequences through a forward noising process and then iteratively reconstructing them with a learned reverse process. This is a key enabler recognized by this work, which exploits this modus operandi to impose global constraints while simultaneously maintaining high fidelity.
@@
where $\alpha_{t}$ is a schedule that decreases over time, $\mathrm{Cat}(\cdot;p)$ is the categorical distribution over probability vector $p\in\Delta^{N}$, and $\nu$ is a fixed reference distribution that specifies the type of corruption applied to each token. For example, the Masked Diffusion Language Model (MDLM) [sahoo2024simple] uses a one-hot vector corresponding to the [MASK] token for $\nu$. In contrast, Uniform Diffusion Language Models (UDLM) [schiff2024simple], uses $\nu$ as the uniform distribution over the vocabulary. Conversely, this instantiation allows tokens to be re-perturbed in later time steps.
@@
$p_{\theta}(\bm{x}_{s}\mid\bm{x}_{t})=q(\bm{x}_{s}\mid\bm{x}_{t},\bm{x}_{0}=\mathbf{x}_{\theta}(\bm{x}_{t},t))\;=\;\begin{cases}\mathrm{Cat}\bigl(\bm{x}_{s};\bm{x}_{t}\bigr),&\text{if }\bm{x}_{t}\neq\bm{\nu},\\[4.0pt] \mathrm{Cat}\!\Bigl(\bm{x}_{s};\frac{(1-\alpha_{s})\,\bm{\nu}+(\alpha_{s}-\alpha_{t})\mathbf{x}_{\theta}(\bm{x}_{t},t)}{1-\alpha_{t}}\,\Bigr),&\text{if }\bm{x}_{t}=\bm{\nu}.\end{cases}$ (2)
Equation (2) formalizes this reverse process parameterization, where $\mathbf{x}_{\theta}(\bm{x}_{t},t)$ denotes the trained denoiser and $s$ is the subsequent timestep in the continuous time reverse process and $0\leq s<t\leq T$. Despite different corruption kernels (absorbing [MASK] [sahoo2024simple] vs. uniform [schiff2024simple]), both samplers expose the full token–probability tensor $x_{t}$ for all positions at every reverse step.
4 Constrained Discrete Diffusion (CDD)
4.1 Projected diffusion sampling.
We begin by introducing a perspective of the discrete diffusion reverse process that motivates our approach. Prior work has shown that continuous score-based diffusion sampling processes can be framed as an sequential optimization procedure [christopher2024constrained]. Score-based parameterizations enable this framing as the model learns to predict the gradients of the data density function, $\mathbf{x_{\theta}}(\bm{x}_{t},t)\approx\nabla\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})$ through Score Matching, and these gradients can then be applied to optimize $\bm{x}_{t}$ to the data distribution $q_{t}(\bm{x}_{t})$. This is typically conducted through a Langevin dynamics algorithms [welling2011bayesian], which is used either directly [song2019generative] or through predictor-corrector Euler-discretizations [song2020score]. In both cases, a series of $M$ steps of Langevin dynamics are taken for a fixed distribution $q_{t}(\bm{x}_{t})$:
@@ 4 Constrained Discrete Diffusion (CDD)
Discrete diffusion models can leverage a discrete generalization of the score function, referred to as the Concrete score, to approximate the gradient of the probability density function [sahoo2024simple, loudiscrete, meng2022concrete]. Concrete Score Matching provides an approach mirroring continuous Score Matching in that the estimated gradients of the probability density function are used to guide the sample to high density regions of the target distribution. While not always explicitly framed as Concrete Score Matching [loudiscrete, meng2022concrete], the denoiser often implicitly models the score function, as supported by theoretical results in [sahoo2024simple] demonstrating its equivalence in simplified formulations ($\nabla\log q_{t}(\bm{x}_{t}\mid\bm{x}_{0})\approx\langle\mathbf{x}_{\theta}(\bm{x}_{t},t),\mathbf{y}\rangle\;\;\forall\mathbf{y}=1,\ldots,N$, in this case). This enables the use of Langevin-based samplers for discrete diffusion, as commonly employed, e.g., in [meng2022concrete].
Effectively, under some regularity conditions, Langevin dynamics will converge towards a stationary point. As shown by xu2018global, Langevin dynamics acts as an “almost-minimizer” on the optimized function, which in this case will be the negative probability density function, converging within a fixed bound of the optimum. Hence, for each series of steps with a stationary density function, within this fixed bound, the sampling procedure can be framed as the optimization problem,
@@
Here, Equation (4) extends the representation from an unconstrained optimization problem to a constrained version by introducing a constraint set $\mathbf{C}$. However, while iteratively applying the denoiser enables sampling from the posterior distribution, it does not natively incorporate externally defined constraints, or even implicit ones, given the stochastic nature of the process. Previous work for continuous modalities has proposed enforcing $\bm{x}_{t}\in\mathbf{C}$ by applying a Euclidean projection after each denoising step [christopher2024constrained], which is natural for continuous modalities which fall in a real space, but this is misaligned when applied to discrete diffusion which operates over a probability simplex $\bm{x}_{t}\in\Delta^{N}$.
@@
Importantly, while state transitions are imposed on the token probability distributions, constraint satisfaction is evaluated on the decoded sequence. Indeed, the constraints are formulated such that the $\arg\max$ of the projected sequence $\bm{y}$ (referred to as $\bm{y}^{\star}$), must satisfy sentence level criteria $\bm{y}^{\star}\in\mathbf{C}$. [^1] We assume a greedy decoding scheme as is standard to current diffusion-based language models. However, this use of a non-differentiable $\arg\max$ operation also poses challenges to the resolution of the projection, which relies on gradient-based optimization.
4.2 Differentiable projection.
To address the non-differentiability of the $\arg\max$ function, we apply a Gumbel-Softmax relaxation $\tilde{\phi}$, which approximates discrete token probabilities with continuous values [jang2017categorical] as
@@
[…]
4.3 Augmented Lagrangian projection.
@@
In practice, $\Delta g$ is non-linear, and, thus, to implement Equation (5), we adopt an augmented Lagrangian approach [boyd2004convex]. In augmented Lagrangian methods, the problem constraints are incorporated into the objective of a minimizer via Lagrange multipliers $\lambda$ and a quadratic penalty term $\mu$. Let $\bm{x}_{t}^{\prime}$ be the probability distribution after the denoising step at diffusion time $t$. We introduce a projected distribution $\bm{y}$, which is iteratively updated to reduce the constraint violations (measured by the score $\Delta g$) while remaining close (in KL-divergence) to $\bm{x}_{t}^{\prime}$. Concretely the augmented Lagrangian dual function is defined as:
@@
and is a lower bound of the original projection operator (5) by weak duality [boyd2004convex]. To obtain the strongest Lagrangian relaxation of the projection, the Lagrangian dual can be used to find the best Lagrangian multipliers $\lambda$ and penalty terms $\mu$, i.e.,
@@
Init: $\lambda,\mu,\eta,\alpha,\delta$;
Input: probability distribution $\bm{x}_{t}^{\prime}$
$\bm{y}\leftarrow\bm{x}_{t}^{\prime}$;
while $\Delta g(\bm{y}^{*})>\delta$ do
 for $j\leftarrow 1$ to max_inner_iter do
    $\mathcal{L}_{KL}\leftarrow D_{\mathrm{KL}}(\bm{x}_{t}^{\prime}\|\bm{y})$
    $\mathcal{L}_{\text{viol}}\leftarrow\sum_{i=1}^{m}\;\lambda_{i}\Delta g_{i}(\tilde{\phi}(\bm{y}))+\;\frac{\mu_{i}}{2}\Delta g_{i}(\tilde{\phi}(\bm{y}))^{2}$
    $\mathcal{L}_{\text{ALM}}\leftarrow\mathcal{L}_{KL}+\mathcal{L}_{\text{viol}}$
    ${\bm{y}}\leftarrow{\bm{y}}-\eta\,\nabla_{\bm{y}}\mathcal{L}_{\text{ALM}}$
 $\lambda\leftarrow\lambda+\mu\,\Delta g(\bm{y}^{*})$
 $\mu\leftarrow\min(\alpha\mu,\,\mu_{\text{max}})$
$\bm{x}_{s}\leftarrow\bm{y}$;
Output: $\bm{x}_{s}$
Algorithm 1 Augmented Lagrangian
The optimization of (7) proceeds iteratively, following a gradient-based update on $\bm{y}$ while dynamically adjusting the Lagrange multiplier $\lambda$ and penalty coefficient $\mu$. Specifically, we perform the following updates [FvHMTBL:ecml20]:
@@
To assess the robustness of this projection formulation, we performed an ablation study (Appendix C) examining the sensitivity of the Lagrangian hyper-parameters. Across the full grid of settings, the Lagrangian relaxation converges to feasible solutions in all our test cases.
@@
Let $\bm{C}$ be non-empty and $\beta$-prox-regular in the sense of [rockafellar2009variational, Def. 13.27], and the score network satisfy $\|\nabla_{\bm{x}_{t}}\log q_{t}(\bm{x}_{t})\|\leq G$ (a standard consequence of the bounded-data domain after normalization). Then, for positive step sizes $\gamma_{t},\leq\frac{1}{2G^{2}}\beta$, the following inequality holds for the distance to the feasible set $\bm{C}$:
@@
Theorem 4.1 shows that the distance to the feasible set $\bm{C}$ decreases at a rate of $1-\bm{\upalpha}_{t}$ at each step (up to an additive $\bm{\upalpha}_{t}^{2}G^{2}$ noise). This implies $\varepsilon$-feasibility after $\mathcal{O}(\bm{\upalpha}_{\min}^{-1})$ steps, with $\bm{\upalpha}_{\min}\propto\min_{t}\gamma_{t}$, and a cumulative KL drift within $\mathcal{O}(\sum_{t}\bm{\upalpha}_{t})$. The theorem assumes $\beta$-prox-regularity, which provides a relaxation of typical convexity assumptions, and implies that for each viable point and normal direction, small perturbations still project uniquely and smoothly back to $\mathbf{C}$. Theorem proofs are provided in Appendix F.